OpenStudy (anonymous):

Differentiate with respect to x: e^(3lnx)

terenzreignz (terenzreignz):

Chain rule.. you know it? :)

OpenStudy (anonymous):

yes, but how??

OpenStudy (anonymous):

my question was, does e^ln cancel out here?

terenzreignz (terenzreignz):

:Whoops I am shamed for not noticing that :D sorry, do the cancellation first

OpenStudy (anonymous):

so what does that become, 3x??

terenzreignz (terenzreignz):

No... remember the properties of logarithms :) \[\huge p \ln x = \ln x^p\]

OpenStudy (anonymous):

or should i do the whole f'(x)e^f(x) thing??

terenzreignz (terenzreignz):

That was what I thought, until you reminded me that e^x and ln x cancel out :) See that exponent \[\huge 3 \ln x\] How can you use that property that I showed you?

OpenStudy (anonymous):

OMG

terenzreignz (terenzreignz):

Stuck? Something wrong?

OpenStudy (anonymous):

I'M AN IDIOT THANK YOU FOR POINTING OUT MY IDIOCY OMG :P

OpenStudy (anonymous):

SO SIMPLE

terenzreignz (terenzreignz):

YOU pointed out my idiocy XD

terenzreignz (terenzreignz):

So what's your answer? :D

OpenStudy (anonymous):

yuck. we suck :P

OpenStudy (anonymous):

3x^2 :D

terenzreignz (terenzreignz):

LOL yeah. Now, for entertainment purposes, let's see what happens if we apply the chain rule... \[\huge \frac{d}{dx}e^{3\ln x}=e^{3lnx}\frac{d}{dx}3\ln x=\frac{3e^{3\ln x}}{x}\] You'd have to simplify the exponential anyway. ANYWAY Great job :)

OpenStudy (anonymous):

omg thats scary looking. Well, it's easier than implicit. Actually, implicit's pretty easy :P

OpenStudy (anonymous):

and omg thank you SO much. stick around, there are some more questions coming up :)

terenzreignz (terenzreignz):

Sure :)

OpenStudy (anonymous):

ok. y = \[x^2\ln (1/x)\]

OpenStudy (anonymous):

where x is the ratio of the inner radius to outer radius. the inner radius is 1/2. for max speed, what should be the outer radius?

OpenStudy (anonymous):

I've got the answer, i just wanna double check. i have \[\sqrt{e}/2\]

terenzreignz (terenzreignz):

I honestly don't know what these mean :) What's an inner radius?

OpenStudy (anonymous):

arrey, imagine a sphere. or say, the earth. so the radius of the core, and the radius of the mantle so to speak.

terenzreignz (terenzreignz):

Okay...